1. Data related to the inventories of Costco Medical Supply are presented below:

Surgical
Equipment
Surgical
Supplies
Rehab
Equipment
Rehab
Supplies
Selling price $ 260 $ 137 $ 333 $ 151
Cost 167 108 280 144
Costs to sell 12 9 16 14


In applying the lower of cost or net realizable value rule, the inventory of rehab equipment would be valued at:

Multiple Choice

  • $267.

  • $280.

  • $317.

  • $234.

2. Data below for the year ended December 31, 2021, relates to Houdini Inc. Houdini started business January 1, 2021, and uses the LIFO retail method to estimate ending inventory.

Cost Retail
Beginning inventory $ 90,000 $ 105,000
Net purchases 416,440 600,000
Net markups 38,000
Net markdowns 58,000
Net sales 537,000


Current period cost-to-retail percentage is:

Multiple Choice

  • 73.9%.

  • 85.7%.

  • 65.3%.

  • 71.8%.
Answer #1
1
Net realizable value = Sales price-Cost to sell
Net realizable value Rehab Equipment = 333-16 = $317
Lower of Cost or Net realizable value for Rehab Equipment is Cost i.e. $280
Option B $280 is correct
2
Cost Retail
Net purchases 416440 600000
Add: Net markups 38000
Less: Net markdowns (58000)
Goods available for sale 416440 580000
Current period cost-to-retail percentage 71.8% =416440/580000
Option D 71.8% is correct